Single Variable Optimization

4. A firm produces Q = 2/L units of a commodity when L units of labor are employed. If the price obtained per unit is $160 and the price per unit of labor is $40, what value of L maximizes profit?
Answer #1

Production function is a function of labor only so optimal quantity of labor hired will be the one where marginal product of labor is equal to real wage

MPL = w/p

2*0.5*(L)^-0.5 = 40/160

L^-0.5 = 0.25

This gives optimal value of L = 0.25^(-1/0.5) = 16.